# The 2010 Census Confidentiality Protections Failed, Here’s How and Why

### v1

@@ 10. The 2020 Census Statistical Disclosure Limitation System Addressed the Failures of the 2010 System
$\text{rSWAPLo}_{b,t}$-CEF 500-999 27,029 14,090 24,060 52.1 89.0
$\text{rSWAPHi}_{b,t}$-CEF 500-999 27,029 11,480 19,130 42.5 70.8
@@
$\text{rSWAPLo}_{b,t}$-CEF 1,000+ 14,043 8,798 12,720 62.7 90.6
$\text{rSWAPHi}_{b,t}$-CEF 1,000+ 14,043 7,678 10,730 54.7 76.4
@@
The rows labeled $\text{rMDF}_{b,t}$ and MDF (light gray highlight) are the analogues of $\text{rHDF}_{b,t}$
@@
reconstruction-abetted reidentification attack to the specially swapped versions of HDF
described in the main text.
Figures 5 and 6 compare the putative reidentification, confirmed reidentification, and reidentification precision rates by census block size when the attacker is COMCRL or $\text{CEF}_{atkr}$, respectively. Their interpretation is identical to the interpretation of Figures 3 and 4. Detailed statistics can be found in Appendix Tables 15, 16, and 17.
Figure 5. Comparison of Putative Reidentification Rates, Confirmed Reidentification Rates, and Reidentification Precision Rates for Alternative Disclosure Limitation Implementations Applied to the 2010 Census Edited File for Attacker COMCRL by Census Block Size
Notes: The denominator used in the first column is the COMCRL Population column in Table 15, which totals to 286,700,000. The denominator in the second column is the COMRCL population in Table 16, and the denominator in the third column is the COMRCL population in Table 17. The denominators in the second and third columns sum to 106,300,000 because only that subset of COMRCL records can be classified as modal and nonmodal from the CEF. See Table 3.
@@
Notes: The denominator used in the first column is the $\text{CEF}_{atkr}$ Population column in Table 15, which totals to 276,000,000. The denominator in the second column is the $\text{CEF}_{atkr}$ population in Table 16, and the denominator in the third column is the $\text{CEF}_{atkr}$ population in Table 17. The denominators in the second and third columns sum to 276,000,000 because all records in $\text{CEF}_{atkr}$ can be classified as modal and nonmodal from the CEF. See Table 3.
The second important feature of Figures 5 and 6 is the very low, nearly uniform by block size, reidentification precision rates for $\text{rMDF}_{b,t}$ and MDF for the nonmodal data-defined persons (middle column). These are the only SDL treatments shown in the figures that accomplish this reduction. As the data in Appendix Table 16 show, the reidentification precision rate for nonmodal persons never exceeds 26.4% and averages 20.1% to 20.4%. An attacker using the methods in this paper would be wrong four times in five guessing the {race, ethnicity} from either the reconstructed MDF, $\text{rMDF}_{b,t}$, or the MDF itself. This is essentially the same as the failure rate on nonmodal persons for the PRG statistical baseline, 16.0% to 16.7%, shown in Appendix Table 13.
Table 10 contains the final piece of evidence that the 2020 DAS effectively countered reconstruction-abetted reidentification attacks as we have modeled them here. For the nonmodal persons living in 0 solution variability blocks who are population unique on {block, sex, agebin} (Panel B), the reidentification precision rates of $\text{rMDF}_{b,t}$ and MDF vary between 19.7% and 21.8%, essentially identical to the precision rates for nonmodal persons in general. Neither the low nor high swap rate experiments deliver precision rates for this vulnerable population below 72.2%. An attacker using $\text{rSWAPHi}_{b,t}$ and our methods would be correct about the population unique nonmodal {race, ethnicity} approximately three times out of four, even though 50% of the housing unit records have been swapped.
The replication archive also contains an ExcelTM (rmdf_bt_0solva_extract.xlsx) workbook with reconstructed MDF records from the same 29 census tracts in which every block has 0 solution variability for $\text{rHDF}_{b,t}$ that are shown in rhdf_bt_0solvar_extract.xlsx. Tags also identify every row that is population unique on {block, sex, agebin} and {block, sex, agebin, race, ethnicity} in the spreadsheet. Inspection of the record-level data in these two workbooks shows how the DAS reduced both putative reidentification and reidentification precision rates.
@@
We have only considered the ability of the 2020 DAS to defend against reconstruction-abetted reidentification attacks; however, both the redistricting and DHC data products were extensively tuned for accuracy. These results are available in other sources [USCB:2023:DHC:DEMO:Metrics, Abowd20222020, cumings-menon:etal:2023] and in the replication archive for this paper. The replication archive also contains the same extensive set of metrics for $\text{rSWAPLo}_{b,t}$ and $\text{rSWAPHi}_{b,t}$. The low swap rate experiment has accuracy comparable to SF1. The high swap rate experiment has accuracy far worse than the production DHC on most statistics.
The other feasible alternative to the 2020 DAS was a variant of the suppression methods used in the 1980 Census. Suppression can defend against an important subset of reconstruction attacks—isolating the records of population uniques; however, successful suppression systems, such as the one used for economic censuses [cox:1995], require both primary and complementary suppression. Appendix D lays out the rules used for the 1980 Census as we adapted them to testing suppression for the 2020 Census. Only primary and whole table suppression were used. Without complementary suppression there is no provable defense against reconstruction of population unique records. We considered only the four primary tables in the P.L. 94-171 Redistricting Data Summary File (P8–P11 in Summary File 1). Appendix Tables 18 and 19 show that the use of just the 1980 Census primary suppression rules would have resulted in zeroing out up to 83.8% of the cells in these tables, suppressing up to 87.7% of the block-level redistricting tables, and suppressing 38.7% of the other block-level tables in the 2010 SF1. Suppression destroys the redistricting use case—missing block-level tables prevent the creation of new voting districts, which have unknown boundaries at the time of data publication. The new districts must have approximately equal populations and be drawn in compliance with Section 2 of the 1965 Voting Rights Act. This limitation of suppression was known in 1980 and constituted one of the main reasons for using swapping in the 1990 Census data [mckenna2018disclosure].
@@ 11. Conclusions
This paper directly addresses the concept at the heart of traditional statistical disclosure limitation: the claim that as long as there is some uncertainty regarding whether or not a particular respondent’s data were used in a particular statistic the agency has provided “plausible deniability.” We establish that the 2010 Census attacker does not need access to the confidential data to know (1) that the record is an exact copy of the confidential HDF (using the 38-age-bin schema) and (2) that the record is a population unique on the quasi-identifiers {block, sex, agebin} used for a classic record-linkage attack. While this is strictly true for “only” 97 million persons, adding more tables and more features can only increase this number. If the reconstructed microdata meet the conditions for a microdata-based record-linkage attack, then such an attack must be defended in the SDL framework. Therefore, the plausible deniability argument turns on whether the 2010 Census aggregation and swapping were adequate. We argued here that they were not. In particular, if the released tabular data have reidentification precision rates that are essentially identical to the original confidential data, then either those data could have been released without aggregation, which the Census Bureau has consistently argued would violate Title 13, or the tabular data are too disclosive, which is what we conclude. The burden of proof now lies with traditional SDL to defend the status quo or accept the consequences of our research.
@@
This paper has also demonstrated that the decision to replace the SDL system used for the 2010 Census was based on sound scientific evidence that those methods would fail if applied to the 2020 Census and/or would lead to unacceptable data accuracy loss. First, we demonstrated that based on only 5 billion of the 150 billion statistics published from the 2010 Census, the confidential microdata from the tabulation input file, the Hundred-percent Detail File, could be recreated with at least 95% accuracy on the schema used for all block-level data. This image of the 2010 HDF, $\text{rHDF}_{b,t}$, fails three contemporaneous confidentiality requirements for the 2010 Census: (i) it is not a sample; (ii) the minimum population in the geographic identifier is 1 person, not 100,000; (iii) the minimum national population for cells in one-dimensional marginals is less than 10,000. Similar criteria are still in place for public-use microdata samples for other Census Bureau household surveys. Second, we demonstrated that the Census Bureau was correct to insist on minimum populations in microdata for both the geographic identifiers and demographic characteristics because for more than 70% of all blocks, the records in $\text{rHDF}_{b,t}$ are provably identical to their counterparts in the HDF using the binned-age schema; hence, there is no confidentiality protection from aggregation over the geographic identifier population. Third, we demonstrated that $\text{rHDF}_{b,t}$ could be used to make high-precision confidentiality-violating inferences: (i) the reidentification precision rate on inferences for nonmodal vulnerable populations is at least 62% and at worst 95% whereas the best precision rate for scientific inferences in vulnerable populations is 32%, and a modal guesser never exceeds 9% precision. Finally, we demonstrated that the differentially private SDL methods applied to the 2020 DHC, the successor to the 2010 SF1, successfully counter this reconstruction-abetted reidentification attack by greatly reducing the putative reidentification rates and by limiting the precision rates on nonmodal vulnerable populations to at most 26.4%. Although such a demonstration does not mean the public 2020 Census data are protected against an arbitrary attack into the indefinite future, it does mean that they are substantially better protected than they would have been had the methods applied to the 2010 Census been used.
@@
The views and opinions expressed in this paper are those of the authors and not the U.S. Census Bureau. Research was conducted under Project ID: P-7502798. John Abowd, Simson Garfinkel, Ramy Tadros, and Lars Vilhuber worked in their personal capacities and without access to any confidential data after leaving the Census Bureau (Abowd, Garfinkel, and Vilhuber) and Galois (Tadros), respectively, to assist in preparing the manuscript for publication. Statistics reported were released under DRB Clearance numbers CBDRB-FY20-DSEP-001, CBDRB-FY22-DSEP-003, CBDRB-FY22-DSEP-004, and CBDRB-FY23-0152.
@@ Appendix A Detailed Tables Referenced in Section 8
Alg. 3) Size (×$\times 10^{-3}$) (×$\times 10^{-3}$) (×$\times 10^{-3}$) (%) (×$\times 10^{-3}$) (×$\times 10^{-3}$) (×$\times 10^{-3}$) (%)
@@
significant digits to conform to disclosure limitation requirements. Population for attacker COMRCL
is the total number of data-defined records in COMRCL that are also in the CEF universe (see
Table 3). Population for attacker $\text{CEF}_{atkr}$ is the total number of data-defined CEF records.
@@
Alg. 3) Size (×$\times 10^{-3}$) (×$\times 10^{-3}$) (×$\times 10^{-3}$) (%) (×$\times 10^{-3}$) (×$\times 10^{-3}$) (×$\times 10^{-3}$) (%)
@@
only data-defined records. Only the 106,300,000 records in COMCRL that match CEF on the
feature set {$pik$, $block$} (see Table 3) can be used to identify a nonmodal person. For each
@@
Alg. 3) Size (×$\times 10^{-3}$) (×$\times 10^{-3}$) (×$\times 10^{-3}$) (%) (×$\times 10^{-3}$) (×$\times 10^{-3}$) (×$\times 10^{-3}$) (%)
@@
only data-defined records. Only the 106,300,000 records in COMCRL that match CEF on the
feature set {$pik$, $block$} (see Table 3) can be used to identify a modal person. For each attacker,

### v2

@@ 9. The 2020 Census Statistical Disclosure Limitation System
$\text{rSWAPLo}_{b,t}$ (5%)-CEF 500-999 27,029 14,090 24,060 52.1 89.0
$\text{rSWAPHi}_{b,t}$ (50%)-CEF 500-999 27,029 11,480 19,130 42.5 70.8
@@
$\text{rSWAPLo}_{b,t}$ (5%)-CEF 1,000+ 14,043 8,798 12,720 62.7 90.6
$\text{rSWAPHi}_{b,t}$ (50%)-CEF 1,000+ 14,043 7,678 10,730 54.7 76.4
@@
The rows labeled $\text{rMDF}_{b,t}$ and MDF (light gray highlight) are the analogs of $\text{rHDF}_{b,t}$
@@
reconstruction-abetted reidentification attack to the specially swapped versions of CEF
described in the text and Appendix C.
Figures 5 and 6 compare the putative reidentification, confirmed reidentification, and reidentification precision rates by census block size when the attacker is COMRCL or $\text{CEF}_{atkr}$, respectively. Their interpretation is identical to the interpretation of Figures 3 and 4. Detailed statistics can be found in Appendix Tables LABEL:tab:reid_blockpop_mdf, LABEL:tab:reid_blockpop_nm_mdf, and LABEL:tab:reid_blockpop_m_mdf.
Figure 5. Comparison of Putative Reidentification Rates, Confirmed Reidentification Rates, and Reidentification Precision Rates for Alternative Disclosure Limitation Implementations Applied to the 2010 Census Edited File for Attacker COMRCL by Census Block Size
Notes: The denominator used in the first column is the COMRCL Population column in Table LABEL:tab:reid_blockpop_mdf, which totals to 286,700,000. The denominator in the second column is the COMRCL population in Table LABEL:tab:reid_blockpop_nm_mdf, and the denominator in the third column is the COMRCL population in Table LABEL:tab:reid_blockpop_m_mdf. The denominators in the second and third columns sum to 106,300,000 because only that subset of COMRCL records can be classified as modal and nonmodal from the CEF. See Table 3.
@@
Notes: The denominator used in the first column is the $\text{CEF}_{atkr}$ Population column in Table LABEL:tab:reid_blockpop_mdf, which totals to 276,000,000. The denominator in the second column is the $\text{CEF}_{atkr}$ population in Table LABEL:tab:reid_blockpop_nm_mdf, and the denominator in the third column is the $\text{CEF}_{atkr}$ population in Table LABEL:tab:reid_blockpop_m_mdf. The denominators in the second and third columns sum to 276,000,000 because all records in $\text{CEF}_{atkr}$ can be classified as modal and nonmodal from the CEF. See Table 3.
The second important feature of Figures 5 and 6 is the very low, nearly uniform by block size, reidentification precision rates for $\text{rMDF}_{b,t}$ and MDF for the nonmodal data-defined persons (middle column). These are the only SDL treatments shown in the figures that accomplish this reduction. As the data in Appendix Table LABEL:tab:reid_blockpop_nm_mdf show, the reidentification precision rate for nonmodal persons never exceeds 26.4% and averages 20.1% to 20.4%. An attacker using the methods in this paper would be wrong four times in five guessing the {race, ethnicity} from either the reconstructed MDF, $\text{rMDF}_{b,t}$, or the MDF itself. This is essentially the same as the failure rate on nonmodal persons for the PRG statistical baseline, 16.0% to 16.7%, shown in Appendix Table LABEL:tab:reid_blockpop_nm.
Table LABEL:tab:nm_solvar contains the final piece of evidence that the 2020 DAS effectively countered reconstruction-abetted reidentification attacks as we have modeled them here. For the nonmodal persons living in zero solution variability blocks who are population unique on {block, sex, agebin} (Panel B), the reidentification precision rates of $\text{rMDF}_{b,t}$ and MDF vary between 21.6% and 21.8%, essentially identical to the precision rates for the proportional guesser (20.9% and 20.2% for COMRCL and $\text{CEF}_{atkr}$, resp.). Neither the low nor high swap rate experiments deliver precision rates for this vulnerable population below 76.4%. An attacker using $\text{rSWAPHi}_{b,t}$ and our methods would be correct about the population unique nonmodal {race, ethnicity} approximately three times out of four, even though 50% of the housing unit records have been swapped.
The replication archive also contains an ExcelTM (rmdf_bt_0solva_extract.xlsx) workbook with reconstructed MDF records from the same 29 census tracts in which every block has zero solution variability for $\text{rHDF}_{b,t}$ that are shown in rhdf_bt_0solvar_extract.xlsx. Tags also identify every row that is population unique on {block, sex, agebin} and {block, sex, agebin, race, ethnicity} in the spreadsheet. Inspection of the record-level data in these two workbooks shows how the DAS reduced both putative reidentification and reidentification precision rates.
@@
We have only considered the ability of the 2020 DAS to defend against reconstruction-abetted reidentification attacks; however, both the redistricting and DHC data products were extensively tuned for accuracy. These results are available in other sources (USCB:2023:DHC:DEMO:Metrics; Abowd20222020; cumings-menon:etal:2023) and in the replication archive for this paper. The replication archive also contains the same extensive set of metrics for $\text{rSWAPLo}_{b,t}$ and $\text{rSWAPHi}_{b,t}$. The low swap rate experiment has accuracy comparable to SF1. The high swap rate experiment has accuracy far worse than the production DHC on most statistics.
The other feasible alternative to the 2020 DAS was a variant of the suppression methods used in the 1980 Census. Suppression can defend against an important subset of reconstruction attacks—isolating the records of population uniques; however, successful suppression systems, such as the one used for economic censuses (cox:1995), require both primary and complementary suppression. Appendix D lays out the rules used for the 1980 Census as we adapted them to testing suppression for the 2020 Census. Only primary and whole table suppression were used. Without complementary suppression there is no provable defense against reconstruction of population unique records. We considered only the four primary tables in the Redistricting Data (P.L. 94-171) Summary File (tables P8–P11 in Summary File 1). Appendix Tables 18 and 19 show that the use of just the 1980 Census primary suppression rules would have resulted in zeroing out up to 83.8% of the cells in these tables, suppressing up to 87.7% of the block-level redistricting tables, and suppressing 38.7% of the other block-level tables in the 2010 SF1. Suppression destroys the redistricting use case—missing block-level tables prevent the creation of new voting districts, which have unknown boundaries at the time of data publication. The new districts must have approximately equal populations and be drawn in compliance with Section 2 of the 1965 Voting Rights Act. This limitation of suppression was known in 1980 and constituted one of the main reasons for using swapping in the 1990 Census data (mckenna:2018:disclosure).
@@ 10. Conclusions
This paper directly addresses the concept at the heart of traditional statistical disclosure limitation: the claim that as long as there is any uncertainty regarding whether or not a particular respondent’s data were used in a particular statistic the agency has provided meaningful confidentiality protection to respondents, so-called “plausible deniability.” We establish that the 2010 Census attacker does not need access to the confidential data to reconstruct a close approximation to the confidential HDF, to know that many records are identical in the reconstructed data and the HDF (using the 38-age-bin schema), and, in many cases, to know which records those are (see Section 6.1). While this is strictly true for “only” 97 million persons, adding more tables and more features can only increase this number. If the reconstructed microdata meet the conditions for a microdata-based record-linkage attack, then such an attack must be defended in the SDL framework. Therefore, the plausible deniability argument turns on whether the 2010 Census aggregation and swapping were adequate. We argued here that they were not. In particular, if the released tabular data have reidentification precision rates that are essentially identical to the original confidential data as in Table LABEL:tab:nm_solvar, then either those data could have been released without aggregation, which the Census Bureau has consistently argued would violate Title 13, or the tabular data are too disclosive, which is what we conclude. From the lens of discussion of on why protections are necessary, any respondent who does not wish their or their child’s race or ethnicity to be learned from the 2010 Census publications may understandably find these results concerning, since they imply that a reconstruction-abetted reidentification attack could infer the correct race and ethnicity with high confidence. Furthermore, such an attack only succeeds because the respondent provided this information (see Section 2.2). Although many other respondents may not find inference about their race and ethnicity concerning, we also remind the reader that race and ethnicity were used here as an example for illustration, not because of any special structure these variables have that is useful for a reconstruction-abetted reidentification attack. Possibly, many more respondents would be concerned if a similar attack could exploit their specific 2010 Census responses to infer whether they were in a same-sex marriage, whether their child was adopted, whether they were an older person living alone, if they exceeded their lease’s occupancy limit, etc.
@@
This paper has also demonstrated that the decision to replace the SDL system used for the 2010 Census was based on sound scientific evidence that those methods would fail if applied to the 2020 Census and/or would lead to unacceptable data accuracy loss. First, we demonstrated that based on only 5 billion of the 150 billion statistics published from the 2010 Census, the confidential microdata from the tabulation input file, the Hundred-percent Detail File, could be recreated with at least 95% accuracy on the schema used for all block-level data (see Section 6.2 and Table LABEL:tab:reconstruction_by_blocksize). This image of the 2010 HDF, $\text{rHDF}_{b,t}$, fails three contemporaneous confidentiality requirements for the 2010 Census: (i) it is not a sample; (ii) the minimum population in the geographic identifier is 1 person, not 100,000; (iii) the minimum national population for cells in one-dimensional marginals is less than 10,000 (see Sections 2 and 8). Similar criteria are still in place for public-use microdata samples for other Census Bureau household surveys. Second, in Table 6, we demonstrated that the Census Bureau was correct to insist on minimum populations in microdata for both the geographic identifiers and demographic characteristics because for more than 70% of all blocks, the records in $\text{rHDF}_{b,t}$ are provably identical to their counterparts in the HDF using the binned-age schema; hence, there is no confidentiality protection from aggregation over the geographic identifier population. Third, we demonstrated that $\text{rHDF}_{b,t}$ could be used to make high-precision confidentiality-violating inferences: as shown in Table LABEL:tab:nm_solvar, the reidentification precision rate on inferences for nonmodal vulnerable populations is at least 60.3% and at worst 95.4% whereas the best precision rate for scientific inferences in vulnerable populations is 20.9%, and a modal guesser never exceeds 3.3% precision. Finally, we demonstrated that the differentially private SDL methods applied to the 2020 DHC, the successor to the 2010 SF1, successfully counter this reconstruction-abetted reidentification attack by greatly reducing the putative reidentification rates and by limiting the precision rates on nonmodal vulnerable populations to at most 21.8% (see Table LABEL:tab:nm_solvar). Although such a demonstration does not mean the public 2020 Census data are protected against an arbitrary attack into the indefinite future, it does mean that they are substantially better protected than they would have been had the methods applied to the 2010 Census been used.
@@
The views and opinions expressed in this paper are those of the authors and not the U.S. Census Bureau. Research was conducted under Project ID: P-7502798. John Abowd, Tamara Adams, Robert Ashmead, Simson Garfinkel, Nathan Goldschlag, Philip Leclerc, Ethan Lew, Ramy Tadros, and Lars Vilhuber worked in their personal capacities and without access to any confidential data after leaving the Census Bureau (Abowd, Adams, Ashmead, Garfinkel, Goldschlag, Leclerc, and Vilhuber) and Galois (Lew and Tadros), respectively, to assist in preparing the manuscript for publication. Statistics reported were released under DRB Clearance numbers CBDRB-FY20-DSEP-001, CBDRB-FY22-DSEP-003, CBDRB-FY22-DSEP-004, and CBDRB-FY23-0152. This version of the full technical report was produced on July 23, 2025. It is the accepted Harvard Data Science Review supplemental text before publication copyediting. This version supersedes previous versions including the one entitled “The 2010 Census Confidentiality Protections Failed, Here’s How and Why” (https://doi.org/10.48550/arXiv.2312.11283, version 1). The published version can be found at https://doi.org/10.1162/99608f92.4a1ebf70.
@@ Appendix A Detailed Tables Referenced in Section 7
Alg. 3) Size (×$\times 10^{3}$) (×$\times 10^{3}$) (×$\times 10^{3}$) (%) (×$\times 10^{3}$) (×$\times 10^{3}$) (×$\times 10^{3}$) (%)
@@
significant digits to conform to disclosure limitation requirements. Population for attacker
COMRCL is the total number of data-defined records in COMRCL that are also in the CEF
universe, per Table 3. Population for attacker $\text{CEF}_{atkr}$ is the total number of data-defined
CEF records.
@@
Alg. 3) Size (×$\times 10^{3}$) (×$\times 10^{3}$) (×$\times 10^{3}$) (%) (×$\times 10^{3}$) (×$\times 10^{3}$) (×$\times 10^{3}$) (%)
@@
only data-defined records. Only the 106,300,000 records in COMRCL that match CEF on the
feature set {$pik$, $block$}, per Table 3, can be used to identify a nonmodal person. For each
@@
Alg. 3) Size (×$\times 10^{3}$) (×$\times 10^{3}$) (×$\times 10^{3}$) (%) (×$\times 10^{3}$) (×$\times 10^{3}$) (×$\times 10^{3}$) (%)
@@
only data-defined records. Only the 106,300,000 records in COMRCL that match CEF on the
feature set {$pik$, $block$}, per Table 3), can be used to identify a modal person. For each attacker,